Circular No. 6535

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997B IN IC 438 Alessandro Gabrijelcic, Passo Valles, Italy, reports his discovery of a

supernova (mag ∼ 16.5) on a CCD image taken on Jan. 13.91 UT. The new object was reported by Gabrijelcic as being 40′′ northeast of the center of IC 438 (α = 5^{h}53^{m}.0, δ = −17^{o}53′, equinox 2000.0); nothing is present at this location on the Digitized Sky Survey.

S. Benetti and C. Lidman, European Southern Observatory (ESO), report: “Inspection of a fully-reduced spectrum (range 523.6–903.1 nm, resolution ∼ 2.0 nm), obtained at Jan. 14.03 UT with the ESO 3.6-m telescope (+ EFOSC1), confirms SN 1997B to be like a type-Ic supernova, ∼ 10 days past maximum. The spectrum is dominated by a huge Ca II infrared line with a P-Cyg profile (expansion velocity ∼ 15 800 km/s) superimposed on an almost-flat continuum. The redshift of the galaxy, 3246 km/s, is determined from the Hα line of a nearby H II region. A broad emission feature (FWHM ∼ 45 nm), showing several structures, is centered at ∼ 661.3 nm — the strongest of which is a resolved emission line (FWHM ∼ 4.6 nm) centered at 663 nm, which could be identified as Hα. The O I 777.3-nm line is also present and relatively strong, with a P-Cyg profile (expansion velocity ∼ 7500 km/s). A P-Cyg line with a flat absorption profile (at ∼ 575.8 nm) is also recognized in the spectrum, and if we identify this feature with He I 587.6-nm, an expansion velocity of ∼ 9200 km/s is deduced. The supernova is located 42′′ east and 11′′.5 north of the galaxy nucleus, and the magnitude deduced from a R-band frame is 16.6.”

EY CYGNI A rare outburst of this dwarf-nova-type (SS-Cyg subclass) cataclysmic

variable EY Cyg has been reported independently by G. Dyck, Assonet, MA; by J. Pietz, Erftstadts, Germany; and by P. Schmeer, Bischmisheim, Germany. E. Waagen, AAVSO, notes that the most recent outbursts of this star, as recorded in the AAVSO International Database, occurred in 1992 Mar. and 1986 Aug. The following visual magnitude estimates are from the AAVSO: 1997 Jan. 10.08 UT, [14.5 (G. Hanson, Cave Creek, AZ); 11.71, 12.8: (Pietz); 11.965, 13.1 (Dyck); 12.097, 12.7 (Hanson); 12.118, 12.7 (Hanson); 12.46, 12.0 (Dyck); 12.715, 12.5: (Schmeer); 12.731, 12.4 (Schmeer); 13.733, 12.1 (Schmeer).

1997 January 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green